31.)Show the calculation of the boiling point of a solution made by dissolving 15.9 grams of the nonelectrolyte C3H7OH in 350 grams of water. Kb for water is 0.51, BP of pure water is 100oC.

30.)Show the calculation of the freezing point of a solution made by dissolving 12.6 grams of the nonelectrolyte C2H5OH in 200 grams of water. Kf for water is 1.86, FP of pure water is 0oC.

29.)Show the calculation of the volume of 0.238 M solution which can be prepared using 13.4 grams of Ca3(PO4)2.

28.)Show the calculation of the volume of 0.987 M solution which can be prepared using 24.6 grams of NaNO3.
Answer #1
  1. To calculate the boiling point elevation, we use the formula:

ΔTb = Kb * molality

First, we need to calculate the molality of the solution:

molality = moles of solute / mass of solvent in kg

We have 15.9 g of C3H7OH, whose molar mass is 60.1 g/mol. Thus, we have:

moles of C3H7OH = 15.9 g / 60.1 g/mol = 0.265 mol

We have 350 g of water, which is 0.350 kg. Thus, we have:

molality = 0.265 mol / 0.350 kg = 0.757 mol/kg

Now we can calculate the boiling point elevation:

ΔTb = 0.51 * 0.757 = 0.386°C

Therefore, the boiling point of the solution is:

Boiling point = 100°C + 0.386°C = 100.386°C

  1. To calculate the volume of solution, we use the formula:

moles of solute = Molarity * volume of solution in liters

First, we need to calculate the moles of NaNO3:

moles of NaNO3 = 24.6 g / 85.0 g/mol = 0.289 mol

Now we can calculate the volume of solution:

volume of solution = moles of solute / Molarity

Molarity = 0.987 mol/L

volume of solution = 0.289 mol / 0.987 mol/L = 0.293 L

Therefore, the volume of solution is 0.293 L or 293 mL.

  1. To calculate the volume of solution, we use the formula:

moles of solute = Molarity * volume of solution in liters

First, we need to calculate the moles of Ca3(PO4)2:

moles of Ca3(PO4)2 = 13.4 g / (340.1 g/mol + 230.97 g/mol + 8*16.0 g/mol) = 0.0348 mol

Now we can calculate the volume of solution:

volume of solution = moles of solute / Molarity

Molarity = 0.238 mol/L

volume of solution = 0.0348 mol / 0.238 mol/L = 0.146 L

Therefore, the volume of solution is 0.146 L or 146 mL.

  1. To calculate the freezing point depression, we use the formula:

ΔTf = Kf * molality

First, we need to calculate the molality of the solution:

molality = moles of solute / mass of solvent in kg

We have 12.6 g of C2H5OH, whose molar mass is 46.1 g/mol. Thus, we have:

moles of C2H5OH = 12.6 g / 46.1 g/mol = 0.273 mol

We have 200 g of water, which is 0.200 kg. Thus, we have:

molality = 0.273 mol / 0.200 kg = 1.365 mol/kg

Now we can calculate the freezing point depression:

ΔTf = 1.86 * 1.365 = 2.54°C

Therefore, the freezing point of the solution is:

Freezing point = 0°C - 2.54°C = -2.54°C


answered by: Hydra Master